Grade upto college level Physical Chemistry

Calculate the molaity of 1 litre solution of 93% H2SO4 (weight/volume). The density of the solution is 1.84 g/ml.

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
Molality = Mass of solute/M. wt. of solute/Mass of solvent in kg
Mass of H­2SO4 in 100 ml of 93% H2SO4­ solution = 93 g
∴ Mass of H2SO­4 in 1000 ml of the H2SO4 solution = 930 g
Mass of 1000 ml H2SO4 solution = 1000 * 1.84 = 1840 g
Mass of water in 1000 ml of solution = 1840 – 930 g
= 910 g = 0.910 kg
Moles of H2SO4 = Wt. of H2SO4/Mol. wt. of H2SO4 = 930/98
∴ Moles of H2SO4 in 1 kg of water
= 930/98 * 1/0.910 = 10.43 mol
∴ Molality of solution = 10.43m